# When Video Coding Meets Multimodal Large Language Models: A Unified Paradigm for Video Coding

### v1

@@ Introduction
IT2V generative model. We perform ablation studies focused on the different settings, including the influence of motion description, different codecs, updating strategies, training step, and sampling step. In terms of motion description, we compare the model without motion description, as depicted in Table 3. The results indicate that incorporating motion description significantly enhances video reconstruction quality. Additionally, we explore a range of codecs for keyframe images, such as Hyperprior (Ballé et al. 2018), NIC (Chen et al. 2021), and NTVP. Among these, NTVP stands out by demonstrating superior reconstructed quality while maintaining a lower coding rate. Our model requires updating $w_{i}$ and $w_{l}$ based on the input, such that we further evaluate the effectiveness of updating strategies, as illustrated in Table 3. To assess the effectiveness of these updating strategies, we present further evaluations in Table 3. Moreover, we examine the repercussions of varying training and sampling steps. An increase in the number of sampling steps correlates with improved results. To strike a balance between performance and computational efficiency, we choose 100 training steps and 50 sampling steps for our final implementation.
@@
Conclusion and Discussion
@@
CMVC includes the TT2V mode and IT2V mode, but more modes can be further explored. For instance, motion representation can be realized through optical flow or trajectories. By integrating multiple modalities of keyframes and motion, we can cater to diverse reconstruction requirements. Furthermore, future efforts should prioritize enhancing CMVC at higher bitrates by integrating more control information to facilitate the reconstruction of the original video. This approach aims to achieve superior performance across all bitrates and dimensions when compared to traditional codecs.

### v2

@@ Introduction
IT2V generative model. We perform ablation studies focused on the different settings, including the influence of motion description, different codecs, updating strategies, training step, and sampling step. In terms of motion description, we compare the model without motion description, as depicted in Table 3. The results indicate that incorporating motion description significantly enhances video reconstruction quality. Additionally, we explore a range of codecs for keyframe images, such as Hyperprior (Ballé et al. 2018), NIC (Chen et al. 2021), and NVTC (Feng et al. 2023). Among these, NVTC stands out by demonstrating superior reconstructed quality while maintaining a lower coding rate. Our model requires updating $w_{i}$ and $w_{l}$ based on the input, such that we further evaluate the effectiveness of updating strategies, as illustrated in Table 3. To assess the effectiveness of these updating strategies, we present further evaluations in Table 3. Moreover, we examine the repercussions of varying training and sampling steps. An increase in the number of sampling steps correlates with improved results. To strike a balance between performance and computational efficiency, we choose 100 training steps and 50 sampling steps for our final implementation.
@@
Discussion
Application. The proposed method enables efficient transmission at low bitrates while preserving semantic content, making it ideal for bandwidth-limited or emergency alert scenarios. When bandwidth allows, keyframe data can be transmitted with textual descriptions, allowing the decoder to reconstruct an approximate video. This hybrid approach balances visual quality with data efficiency, suitable for situations where full video streaming is infeasible, such as disaster response and remote surveillance. Although still in the research phase, the method shows strong potential for real-world applications. With advances in computational resources and model optimization techniques like pruning and quantization, it is expected to become a practical solution for emergency communications and other bandwidth-constrained environments.
Higher bitrate. CMVC includes the TT2V mode and IT2V mode, but more modes can be further explored. For instance, motion representation can be realized through optical flow or trajectories. By integrating multiple modalities of keyframes and motion, we can cater to diverse reconstruction requirements. Furthermore, future efforts should prioritize enhancing CMVC at higher bitrates by integrating more control information to facilitate the reconstruction of the original video. This approach aims to achieve superior performance across all bitrates and dimensions when compared to traditional codecs.
Conclusion